Dad is three times as old as his son. In nine years dad will be twice as old as his son how old is each now?
Answer by fractalier(6550)

Call Dad's and son's age NOW, d and s. Thus
d = 3s
In nine years,
d + 9 = 2(s + 9)
Now substitute in for d here and get
3s + 9 = 2(s + 9)
3s + 9 = 2s + 18
and
s = 9 so that Dad is
d = 27